OpenStudy (rainbow_dash):

Find the inverse of the quadratic function: y=x^2+6x+9

OpenStudy (anonymous):

y=x^2+6x+9 \[y = \left( x+3 \right)^{2}\]

OpenStudy (anonymous):

It's quadratic so it won't be injective. You'll need to split it where the derivative is zero, and then choose one of the sides to take the inverse on. So the positive or negative of the square root that telltoamit derived.

OpenStudy (anonymous):

I mean you want to solve for x above.... so you'll need to take square root, and then choose either the positive or negative of the square root since they're both valid, and that'll correspond to splitting it vertically where the derivative is zero.

OpenStudy (anonymous):